meliahmarie01:

Shown below are the steps a student took to solve the simple interest formula A = P (1 + rt) for r. A = P (1 + rt) A = P + rt A - P = rt r = (A - P)/t Explain the algebraic error in the students work. Then state how the equation r = (A - P)/t can be modified to produce the correct result.

mhchen:

A = P (1 + rt) A = P + rt Take a closer look at this step

meliahmarie01:

Okay?

mhchen:

Do you see the problem with this step?

TheSmartOne:

Let's try to solve it ourselves and then we can compare to see where they went wrong. A = P(1 + rt) We're trying to solve for 'r' What's the first step we would do? We would want to distribute the P into the parentheses Are you aware of the distributive property? a(b+c) = ab + ac You have to multiple the coefficient to every term inside the parentheses. So what do you get when you distribute the P? A = P(1+rt) A = ??